What is the equation for this word problem. The price you pay for a box of donuts varies linearly with the number of donuts in the box. For five donuts the price is 1.15, and for eleven the cost is 2.35. Can you please help me!
Answer by richwmiller(17219)

the two points are (11,2.35) and (5,1.15)
use (y-y)/(x-x)=m
find m
then plug the new m into
1.15=m(5)+b
finally plug in your new m and new b into y=mx+b